# SIMPLE: Simultaneous Multi-Plane Self-Supervised Learning for Isotropic MRI Restoration from Anisotropic Data

### v1

@@ Abstract
Magnetic resonance imaging (MRI) is crucial in diagnosing various abdominal conditions and anomalies. Traditional MRI scans often yield anisotropic data due to technical constraints, resulting in varying resolutions across spatial dimensions, which limits diagnostic accuracy and volumetric analysis. Super-resolution (SR) techniques aim to address these limitations by reconstructing isotropic high-resolution images from anisotropic data. However, current SR methods often rely on indirect mappings and limited training data, focusing mainly on two-dimensional improvements rather than achieving true three-dimensional isotropy. We introduce “SIMPLE,” a Simultaneous Multi-Plane Self-Supervised Learning approach for isotropic MRI restoration from anisotropic data. Our method leverages existing anisotropic clinical data acquired in different planes, bypassing the need for simulated downsampling processes. By considering the inherent three-dimensional nature of MRI data, SIMPLE ensures realistic isotropic data generation rather than solely improving through-plane slices. This approach’s flexibility allows it to be extended to multiple contrast types and acquisition methods commonly used in clinical settings. Our experiments show that SIMPLE outperforms state-of-the-art methods both quantitatively using the Kernel Inception Distance (KID) and semi-quantitatively through radiologist evaluations. The generated isotropic volume facilitates more accurate volumetric analysis and 3D reconstructions, promising significant improvements in clinical diagnostic capabilities. Our code base will be available upon acceptance.
@@ 1 Introduction
Magnetic resonance imaging (MRI) is a cornerstone in medical diagnostics, providing high-resolution images of soft tissues with unparalleled contrast. In abdominal imaging, MRI plays a crucial role in diagnosing and monitoring various abdominal conditions and anomalies such as liver cirrhosis, pancreatic tumors, and renal anomalies [5, 9, 18, 10, 7]. However, traditional MRI scans often produce anisotropic data, where the resolution differs across spatial dimensions, potentially limiting diagnostic accuracy and the ability to perform precise volumetric analysis [4, 21].
@@
The anisotropy in MRI data arises primarily from the technical constraints and trade-offs inherent in the imaging process. Achieving high spatial resolution in MRI requires long acquisition times due to the need to sample a large number of spatial frequencies in k-space, the Fourier domain representation of the image [1]. In clinical settings, the time available for MRI scans is limited by patient comfort, motion artifacts, and the need for high throughput. In abdominal imaging, these constraints are further compounded by the challenges of capturing high-quality images of a region that is prone to motion artifacts from respiration and peristalsis. As a result, compromises are often made, leading to higher resolution in the in-plane dimensions (x and y) and lower resolution in the through-plane dimension (z).
@@
To mitigate the effects of anisotropy and improve diagnostic accuracy, clinical practice often involves acquiring at least two volumes with high in-plane resolution, typically in axial and coronal planes [8]. This approach ensures that although one dimension may have a lower resolution in each individual volume, combining multiple orientations provides a more comprehensive and detailed view of anatomical structures. However, this method can still result in partial volume effects and diminished image quality, particularly in the slice-selection direction, leading to suboptimal volumetric analysis and 3D reconstructions. Furthermore, acquiring two planes increases the total scan time.
@@
Super-resolution (SR) techniques have emerged as powerful tools to address these limitations, enabling the reconstruction of isotropic high-resolution images from anisotropic data [12]. These methods leverage advanced computational algorithms, often using neural networks, to learn a mapping from low-resolution (LR, with large inter-slice spacing) images to high-resolution (HR, with small inter-slice spacing) images. By enhancing the spatial resolution of MRI images, SR techniques facilitate better visualization of anatomical structures and pathological conditions, improve diagnostic performance, aid in surgical planning, and enhance the accuracy of quantitative assessments.
@@
Figure 2: Model Architecture: The dashed red volume denotes the linearly interpolated volume. The blue-framed images represent the ATME-generated single-plane high-quality coronal slices, while the green-framed images represent the ATME-generated single-plane high-quality axial slices. The red dashed arrows indicate sampling along the coronal plane, and the blue dashed arrows indicate sampling along the axial plane. Gray dashed arrows depict the ATME flow, and green dashed arrows highlight the connections to consistency loss.
@@
However, the abdomen poses unique challenges for MRI due to motion artifacts from respiration and peristalsis, and the need for high contrast between adjacent structures. Consequently, acquiring isotropic data for training purposes is not always feasible, particularly in clinical abdominal imaging. This limitation underscores the necessity for SR methods that can effectively work with the anisotropic data typically available in clinical settings.
@@
In an effort to circumvent the need for real high-resolution (HR) data for supervision, recent studies developed self-supervised methodologies, classified into resampling-based and synthesis-based techniques. Resampling-based techniques initially simulate images of even lower resolution from the original low-resolution (LR) images, establish a mapping from these lower-resolution images to the LR images, and subsequently apply this mapping to the LR images to forecast the HR images [13]. Synthesis-based techniques, on the other hand, synthesize HR images from the original LR images and train the super-resolution (SR) model on these synthesized images [27, 20, 16].
Despite their innovation, current self-supervised SR methodologies are hindered by limited training data and indirect mapping strategies. Notably, the training sets are derived from a limited number of LR cases, and the SR mapping is learned from lower-resolution to low-resolution data or from synthesized LR to synthesized HR data, as opposed to actual LR-HR pairs. This indirect mapping could significantly impede the performance of the model. Moreover, existing approaches primarily focus on interpolating the missing slices in the anisotropic volume, rather than simultaneously accounting for the inherent three-dimensional nature of the data.
In this work, we propose “SIMPLE,” a Simultaneous Multi-Plane Self-Supervised Learning approach for Isotropic MRI Restoration from Anisotropic Data. Our method leverages existing anisotropic clinical data acquired in different planes, avoiding the reliance on approximated simulated downsampling processes. Additionally, our approach considers the inherent three-dimensional nature of the data, rather than solely focusing on improving the quality of the two-dimensional slices in the through-plane direction. The flexibility of our approach in not requiring specialized data allows for seamless extension to multiple contrast types and acquisition methods commonly used in clinical practice. Figure 1 provides a comprehensive overview of our model architecture, including essential pre-processing steps such as linear interpolation and patch extraction. It also highlights the core components: the generator and two discriminators, which are crucial for ensuring simultaneous multi-plane realistic generation of the isotropic MRI volume.
Upon training, our approach efficiently produces a 3D isotropic MRI volume from a single anisotropic scan, alleviating issues related to inevitable abdominal motion between volumes and reducing the overall acquisition time required to obtain anisotropic volumes in two distinct planes. The generated isotropic volume ultimately facilitates more accurate volumetric analysis and 3D reconstructions.
@@
Refer to captionAxial VolumeCoronal VolumeAxial ATMECoronal ATMESMORE4SIMPLE Axial View Coronal View Sagittal View
@@
Figure 3: Multi-plane view of slices sampled from the isotropic MRI volume generated by SIMPLE in comparison to five competitive methods (from left to right): linear interpolation on anisotropic axial volume, linear interpolation on anisotropic coronal volume, ATME on anisotropic axial volume, ATME on anisotropic coronal volume, SMORE4 on anisotropic coronal volume, and SIMPLE on anisotropic coronal volume. Our approach, SIMPLE, produced higher-quality slices simultaneously on all planes compared to the other methods, which produced high-quality slices along the super-resolution plane but resulted in artifacts in the other planes.
@@ 2 Previous Work
The application of deep learning for through-plane super-resolution in MRI, which involves learning the non-linear mapping between low and high-resolution images, has seen significant advancements in recent years. Early supervised learning approaches necessitated high-resolution training examples [17, 19, 25]. However, in clinical practice, acquiring such images is often impractical or even impossible, especially in the context of abdominal imaging. To address this limitation, several unsupervised methods have been proposed. Jog et al. [13] proposed an approach where the LR plane is first downsampled further, creating an LR2 dataset. This dataset is then used to build LR-LR2 image pairs, which are employed to learn the mapping from LR2 to LR. Finally, this mapping is applied to the LR data to enhance its resolution.
@@
Zhao et al.[27] used the high-resolution (HR) plane of the anisotropic volume to create HR-low-resolution (LR) image pairs by downsampling the HR plane for their “SMORE” model. The trained model is then applied to the LR plane to enhance its resolution. Remedios et al.[20] improved this by learning the point-spread-function (PSF) in the LR plane to more accurately simulate the downsampling process in the HR plane, creating HR-LR image pairs for training. The trained model is then applied to two LR planes, with the restored volumes averaged into a single HR isotropic volume. Similarly, Lu et al. [16] proposed a Two-stage Self-supervised Cycle-consistency Network (TSCNet) for MR slice interpolation. This method uses a two-stage self-supervised learning (SSL) strategy. In the first stage, paired LR-HR images are synthesized along the sagittal and coronal directions of input LR images for network pretraining. In the second stage, a cyclic interpolation procedure based on triplet axial slices is designed for further refinement.
@@
Zhang et al.[26] used a similar approach to construct HR-LR image pairs for training and adapted the implicit neural representation (INR) network to implement a 2D arbitrary-scale super-resolution (SR) model. Recently, Fang et al.[6] introduced ”CycleINR,” an enhanced INR model for 3D volumetric super-resolution in medical data. They utilized the continuity of the implicit function for super-resolution at arbitrary up-sampling rates, eliminating the need for separate training. Additionally, they improved grid sampling with a local attention mechanism and reduced over-smoothing by incorporating a cycle-consistent loss.
@@
Sander et al. [22] proposed an unsupervised deep learning semantic interpolation approach that synthesizes new intermediate slices from encoded low-resolution examples. They achieved semantically smooth interpolation in the through-plane direction by exploiting the latent space generated by autoencoders. New intermediate slices were generated by combining latent space encodings of two spatially adjacent slices using their convex combination. The combined encoding was then decoded into an intermediate slice.
@@
However, these methods assume a high-quality approximation of the downsampling process and focus on generating high-resolution images in the through-plane, neglecting the three-dimensional isotropic nature of the data. In contrast, our approach uses existing anisotropic clinical data acquired in different planes, avoiding reliance on simulated downsampling. Our method accounts for the inherent 3D nature of the data, rather than solely improving the quality of 2D slices in the through-plane direction.
@@ 3 Method
To address the limitations of previous works and enable simultaneous multi-plane processing, we introduce SIMPLE—a simultaneous multi-plane self-supervised learning model for isotropic super-resolution MRI volume reconstruction from anisotropic MRI data acquired using 2D MRI techniques. Our model effectively manages slice spacing and motion, enhancing the resolution and quality of the 3D MRI volume. This improvement ensures that sampling along any plane (either axial or coronal) results in high-quality 2D slices in both planes. SIMPLE takes as input an anisotropic volume, and produces a low-quality isotropic volume using a linear interpolation. Then, the model restores a high-quality isotropic volume from the low-quality one. Equation 1 provides a mathematical framework for the proposed model. $V_{Cor}$ denotes the anisotropic coronal volume, $V^{\prime}_{Iso}$ represents the linearly interpolated low-quality isotropic volume, $\hat{V}_{Iso}$ signifies the output high-quality isotropic volume, $GM$ denotes the generative model and $L$ the linear interpolation operator. This formulation distinguishes our approach by addressing both constraints, unlike previous methods which typically focus on resolving only one constraint.
$\hat{V}_{\text{Iso}}=GM(V^{\prime}_{\text{Iso}})=GM(L(V_{\text{Cor}}))$ (1)
$s.t:$
@@
$i\in[0,512)$
We implemented our approach by combining a 3D generator based on the U-Net architecture [3] with two 2D conditional patch discriminators [11], specifically designed for the coronal and axial planes. Figure 2 depicts the overall architecture.
@@ 3.1.1 Pre-Processing
As a pre-processing step, we isotropically sample the anisotropic input volume by performing linear interpolation. Due to the significant spacing between the MRI slices, linear interpolation struggles to accurately predict the voxels between the slices, resulting in a low-quality 3D volume. This is reflected in the poor quality of slices in all main 2D planes. To manage the high computational demands of processing the uniform isotropic volume, we extracted 3D patches of size 64$\times$64$\times$64 with an 8% overlap in each dimension.
@@ 3.1.2 Networks
We use two discriminators: one for the coronal plane and another for the axial plane. Unlike the 3D generator, these discriminators operate in the 2D domain. We extract 3D input and output patches along both planes from the generator, creating 64 2D patches of size 64$\times$64 from each 3D patch. Inspired by the pix2pix model [11], each plane has a conditional patch discriminator that takes the input image and either the generated or real image as a single 2-channel input, ensuring the generated image remains realistic and consistent.
@@
Both discriminators use 2D convolutional layers (kernel size 4, stride 2), followed by instance normalization and LeakyReLU activation. Since the volume’s original orientation is in the coronal plane, the generator’s task is simpler in this orientation than in the axial plane. To increase complexity in the coronal plane, we use a simpler architecture for the coronal discriminator, with fewer layers and filters.
@@ 3.1.3 Loss Functions
For the SIMPLE generator, we employ a loss function that equally integrates both coronal and axial components. Each component comprises two types of loss: adversarial loss $\mathcal{L}_{\text{G}_{\text{ADV}}}$, derived from the discriminator’s output classification matrix, and consistency loss $\mathcal{L}_{L_{1}}$, calculated as the L1 distance. For the discriminators, we use only the adversarial loss. The adversarial loss is based on a least square GAN (LSGAN), which provides more stable training compared to vanilla GANs, and conditional patch discriminator. Equation 2 represents the LSGAN loss for the generator, where $x$ denotes the input image.
@@
$\small\begin{split}&\mathcal{L}_{\text{G}}=\alpha\left[\mathcal{L}_{\text{G}_{\text{ADV}_{\text{cor}}}}+\lambda_{\text{cor}}\mathcal{L}_{L_{1_{\text{cor}}}}\right]+\beta\left[\mathcal{L}_{\text{G}_{\text{ADV}_{\text{ax}}}}+\lambda_{\text{ax}}\mathcal{L}_{L_{1_{\text{ax}}}}\right]\end{split}$ (5)
where $\alpha$ and $\beta$ denote the weights of the coronal and axial components, respectively. We set $\alpha$ and $\beta$ both to 0.5.
Refer to captioncase 1case 2case 3(a)(b)(c)(d)
Figure 4: Straight multi-planar reconstructions (MPR) of the terminal ileum in Crohn’s disease patients for three cases based on coronal volumes: (a) anisotropic volume, (b) linearly interpolated isotropic volume, (c) SMORE4 isotropic volume, and (d) SIMPLE isotropic volume. The MPRs generated from the SIMPLE reconstructed volume were less prone to artifacts stemming from the anisotropic MRI acquisition.

### v2

@@ Abstract
Magnetic resonance imaging (MRI) is crucial in diagnosing various abdominal conditions and anomalies. Traditional MRI scans often yield anisotropic data due to technical constraints, resulting in varying resolutions across spatial dimensions, which limits diagnostic accuracy and volumetric analysis. Super-resolution (SR) techniques aim to address these limitations by reconstructing isotropic high-resolution images from anisotropic data. However, current SR methods often depend on indirect mappings and scarce 3D isotropic data for training, primarily focusing on two-dimensional enhancements rather than achieving genuine three-dimensional isotropy. We introduce “SIMPLE,” a Simultaneous Multi-Plane Self-Supervised Learning approach for isotropic MRI restoration from anisotropic data. Our method leverages existing anisotropic clinical data acquired in different planes, bypassing the need for simulated downsampling processes. By considering the inherent three-dimensional nature of MRI data, SIMPLE ensures realistic isotropic data generation rather than solely improving through-plane slices. This approach’s flexibility allows it to be extended to multiple contrast types and acquisition methods commonly used in clinical settings. Our experiments on two distinct datasets (brain and abdomen) show that SIMPLE outperforms state-of-the-art methods both quantitatively using the Kernel Inception Distance (KID), semi-quantitatively through radiologist evaluations, and qualitatively through Fourier domain analysis. The generated isotropic volume facilitates more accurate volumetric analysis and 3D reconstructions, promising significant improvements in clinical diagnostic capabilities. Our code base will be available upon acceptance.
1 Introduction
@@
Magnetic resonance imaging (MRI) is essential in medical diagnostics, offering high-resolution, high-contrast images of soft tissues. In abdominal imaging, MRI is crucial for diagnosing and monitoring conditions like liver cirrhosis, pancreatic tumors, and renal anomalies [6, 10, 20, 11, 8]. However, traditional MRI often yields anisotropic data with differing resolutions across spatial dimensions, limiting diagnostic accuracy and precise volumetric analysis [5, 23].
@@
This anisotropy stems from technical constraints in MRI, where achieving high spatial resolution requires lengthy acquisition times due to extensive “k-space” sampling [2]. In clinical practice, time is limited by patient comfort, potential motion artifacts, and high throughput demands. In abdominal imaging, these challenges are intensified by motion from respiration and peristalsis, leading to compromises with higher in-plane (x, y) resolution and lower through-plane (z) resolution.
@@
To reduce anisotropy and improve diagnostic accuracy, clinicians often acquire at least two high-resolution volumes in different planes, typically axial and coronal [9]. This multi-plane approach provides a more complete view, but partial volume effects and lower quality in the slice-selection direction still limit volumetric analysis and 3D reconstructions. Additionally, acquiring two planes extends scan time.
@@
Super-resolution (SR) techniques offer a solution by reconstructing isotropic high-resolution images from anisotropic data [13]. Using advanced algorithms, often neural networks, SR methods learn to map low-resolution (large inter-slice spacing) images to high-resolution (small inter-slice spacing) images, enhancing anatomical detail, diagnostic accuracy, surgical planning, and quantitative assessments.
@@
Figure 2: Model Architecture: The dashed red volume denotes the linearly interpolated volume. The blue-, green-, and orange-framed images represent the ATME-generated high-quality coronal, axial, and sagittal slices, respectively. Red, blue, and yellow dashed arrows indicate sampling along the coronal, axial, and sagittal planes, respectively. Gray dashed arrows show the ATME flow and green dashed arrows highlight the connections to consistency loss.
@@
The abdomen presents unique MRI challenges due to motion artifacts from respiration and peristalsis, as well as the need for high contrast between structures. This makes acquiring isotropic data for training impractical, especially in clinical abdominal imaging, highlighting the need for SR methods that can handle anisotropic data.
@@
To bypass the requirement for high-resolution (HR) training data, recent studies have developed self-supervised SR methods, categorized into resampling-based and synthesis-based techniques. Resampling-based methods simulate even lower resolution images from low-resolution (LR) data, map these back to the original LR images, and then use this mapping to predict HR images [14]. In contrast, synthesis-based methods generate HR images from LR data and train the SR model on these synthesized images [31, 22, 18].
Current self-supervised SR methods face limitations due to small training datasets and indirect mappings. Typically, these methods rely on a limited number of LR cases, learning SR mappings from either lower-resolution to low-resolution data or synthesized LR to synthesized HR, rather than real LR-HR pairs, which can impact model performance. Moreover, they generally focus on interpolating missing slices without fully considering the 3D structure of the data.
In this work, we introduce “SIMPLE,” a Simultaneous Multi-Plane Self-Supervised Learning approach for Isotropic MRI Restoration from Anisotropic Data. SIMPLE leverages anisotropic clinical scans from multiple planes and addresses the 3D nature of the data, beyond just improving through-plane slice quality. This flexible approach, without requiring specialized data, supports various contrast types and acquisition methods common in clinical practice. Figure 1 illustrates our model architecture.
@@
Refer to captionAxial VolumeCoronal VolumeAxial ATMECoronal ATMEAvg. ATMESMORE4SIMPLE Axial View Coronal View Sagittal View
@@
Figure 3: Multi-plane view of slices sampled from the isotropic MRI volume generated by SIMPLE in comparison to five competitive methods (from left to right): linear interpolation on anisotropic axial volume, linear interpolation on anisotropic coronal volume, ATME on anisotropic axial volume, ATME on anisotropic coronal volume, averaged ATME on coronal and axial planes of anisotropic coronal volume, SMORE4 on anisotropic coronal volume, and SIMPLE on anisotropic coronal volume. SIMPLE produced higher-quality slices simultaneously on all planes compared to the other methods, which produced high-quality slices along the super-resolution plane.
@@ 2 Previous Work
The use of deep learning for through-plane super-resolution in MRI, which learns the non-linear mapping between low- and high-resolution images, has advanced significantly. Early supervised methods required high-resolution training data [19, 21, 28], but these are often impractical in clinical settings, particularly for abdominal imaging. To overcome this, unsupervised methods have been developed. For example, Jog et al. [14] introduced a technique that downscales the LR plane further to create an LR2 dataset, forming LR-LR2 pairs to learn the mapping from LR2 to LR, which is then applied to improve LR resolution.
Slice-to-volume (SVR) [27] and patch-to-volume (PVR) [1] methods integrate axial, coronal, and sagittal images into an isotropic 3D volume. Yucheng et al. [17] proposed a framework for prostate MRI that applies PVR/SVR reconstruction to generate an initial isotropic volume, then re-slices it into 2D slices for SR training across all planes. The three SR volumes are then averaged, yielding a final 3D SR volume with enhanced isotropic resolution.
@@
Zhao et al. [31] used the HR plane in anisotropic volumes to create HR-LR pairs by downsampling for their “SMORE” model, which is then applied to enhance the LR plane. Remedios et al. [22] improved this by learning the point-spread-function (PSF) in the LR plane for more accurate downsampling simulation, creating HR-LR pairs and averaging restored volumes from two LR planes to produce a single HR isotropic volume. Similarly, Lu et al. [18] introduced TSCNet, a two-stage self-supervised learning method that pretrains on synthesized LR-HR pairs in sagittal and coronal planes, followed by cyclic interpolation using triplet axial slices for refinement.
@@
Zhang et al. [30] constructed HR-LR image pairs for training and adapted an implicit neural representation (INR) network for a 2D arbitrary-scale super-resolution (SR) model. Fang et al. [7] later introduced “CycleINR,” an enhanced 3D volumetric SR model that leverages the continuity of implicit functions for arbitrary up-sampling rates without separate training. They further improved performance with local attention for grid sampling and reduced over-smoothing through cycle-consistent loss.
@@
Sander et al. [24] proposed an unsupervised deep learning approach for semantic interpolation, synthesizing intermediate slices from low-resolution inputs. By leveraging the latent space of autoencoders, they achieved smooth through-plane interpolation, generating new slices by convexly combining latent encodings of adjacent slices and decoding the result into an intermediate slice.
@@
However, these methods assume a high-quality approximation of the downsampling process and focus on generating high-resolution images in the through-plane, neglecting the three-dimensional isotropic nature of the data. In contrast, our approach uses existing multi-plane anisotropic clinical data, bypassing simulated downsampling and enhancing quality across all planes to capture the full 3D structure.
@@ 3 Method
To address the limitations of previous works and enable simultaneous multi-plane processing, we introduce SIMPLE—a simultaneous multi-plane self-supervised learning model for isotropic super-resolution MRI volume reconstruction from anisotropic MRI data acquired using 2D MRI techniques. Our model effectively manages slice spacing and motion, enhancing the resolution and quality of the 3D MRI volume. This improvement ensures that sampling along any plane (axial, coronal, or sagittal) results in high-quality 2D slices in all planes. SIMPLE takes as input an anisotropic volume, and produces a low-quality isotropic volume using a linear interpolation. Then, the model restores a high-quality isotropic volume from the low-quality one. Equation 1 provides a mathematical framework for the proposed model. $V_{\text{An-Iso}}$ denotes the anisotropic volume, $V^{\prime}_{\text{Iso}}$ represents the linearly interpolated low-quality isotropic volume, $\hat{V}_{\text{Iso}}$ signifies the output high-quality isotropic volume, $GM$ denotes the generative model and $L$ the linear interpolation operator. This formulation distinguishes our approach by addressing all three constraints, unlike previous methods which typically focus on resolving only one or two.
$\hat{V}_{\text{Iso}}=GM(V^{\prime}_{\text{Iso}})=GM(L(V_{\text{An-Iso}}))$ (1)
@@
[…]
We implemented our approach by combining a 3D generator based on the U-Net architecture [4] with 2D conditional patch discriminators [12], specifically designed for the coronal, axial and sagittal planes. Figure 2 depicts the overall architecture.
@@ 3.1.1 Pre-Processing
As a pre-processing step, we isotropically sample the anisotropic input volume by performing linear interpolation. Due to the significant spacing between the MRI slices, linear interpolation struggles to accurately predict the voxels between the slices, resulting in a low-quality 3D volume. This is reflected in the poor quality of slices in all main 2D planes. To manage the high computational demands of processing the uniform isotropic volume, we extracted 3D patches of size 64$\times$64$\times$64 with an overlap in each dimension.
@@ 3.1.2 Networks
We use two or three discriminators, one for each plane. The number of discriminators depends on data availability: three discriminators are used when the dataset includes anisotropic images from all three planes, while only two discriminators are used if the data includes images from only two planes. Unlike the 3D generator, these discriminators operate in the 2D domain. We extract 3D input and output patches along all planes from the generator, creating 64 2D patches of size 64$\times$64 from each 3D patch. Inspired by the pix2pix model [12], each plane has a conditional patch discriminator that takes the input image and either the generated or real image as a single 2-channel input, ensuring the generated image remains realistic and consistent.
@@
All discriminators use 2D convolutional layers (kernel size 4, stride 2), followed by instance normalization and LeakyReLU activation. Since the volume’s original orientation is in a specific plane, the generator’s task is simpler in this plane compared to others. To increase complexity in the original plane, we use a simpler architecture for the corresponding discriminator, with fewer layers and filters.
@@ 3.1.3 Loss Functions
For the SIMPLE generator, we employ a loss function that equally integrates coronal, axial and sagittal components. Each component comprises two types of loss: adversarial loss $\mathcal{L}_{\text{G}_{\text{ADV}}}$, derived from the discriminator’s output classification matrix, and consistency loss $\mathcal{L}_{L_{1}}$, calculated as the L1 distance. For the discriminators, we use only the adversarial loss. The adversarial loss is based on a least square GAN (LSGAN), which provides more stable training compared to vanilla GANs, and conditional patch discriminator. Equation 2 represents the LSGAN loss for the generator, where $x$ denotes the input image.
@@
$\small\begin{split}\mathcal{L}_{\text{G}}=&\ \alpha\left[\mathcal{L}_{\text{G}_{\text{ADV}_{\text{cor}}}}+\lambda_{\text{cor}}\mathcal{L}_{L_{1_{\text{cor}}}}\right]+\beta\left[\mathcal{L}_{\text{G}_{\text{ADV}_{\text{ax}}}}+\lambda_{\text{ax}}\mathcal{L}_{L_{1_{\text{ax}}}}\right]\\ &\ \ \ \ \ \ \ \ \ \ \ +\gamma\left[\mathcal{L}_{\text{G}_{\text{ADV}_{\text{sag}}}}+\lambda_{\text{sag}}\mathcal{L}_{L_{1_{\text{sag}}}}\right]\end{split}$ (5)
where $\alpha$, $\beta$, $\gamma$ denote the weights of the coronal, axial and sagittal components, respectively. We set $\alpha$, $\beta$ and $\gamma$ both to 0.5.